# Multiplayer Federated Learning: Reaching Equilibrium with Less Communication

### v1

@@ C.2 Quadratic $n$-player game
C.3 Distributed mobile robot control
We follow the same choice of parameter values $a_{i},b_{i},x^{i}_{\mathrm{anc}},h_{ij}$ within (7) from [49]: $n=5$, $d=1$, $a_{i}=10+i/6$, $b_{i}=i/6$,
@@ C.3 Distributed mobile robot control
$\left(h_{ij}\right)_{\begin{subarray}{c}1\leq i\leq 5\\ 1\leq j\leq 5\end{subarray}}=\begin{pmatrix}0&5&-7&9&-8\\ -5&0&-6&2&-9\\ 7&6&0&7&-4\\ -9&-2&-7&0&-2\\ 8&9&4&2&0\end{pmatrix}.$
We add Gaussian noise with $\sigma^{2}=100$ to the gradients to simulate stochasticity. In this setup, all our theoretical assumptions are satisfied.
Appendix D Discussion on Theoretical Assumptions
D.1 Possible simplification of assumptions: Assuming cocoercivity of
@@
In the subsequent paragraphs, we explain in detail why this is the case. However, we emphasize here that if we derived all convergence theory using (COCO) in place of (CVX), (SM) and (SCO) and did not distinguish the role of $L_{i}$’s (the local Lipschitzness parameters from (SM)) from that of $\ell$, then the resulting convergence rates would have become much more pessimistic (worse) in many cases. Therefore, in our work, we choose to use the current set of assumptions. It allows us to more clearly present the tight dependency of convergence rates to $L_{i}$’s. Also note that assuming (CVX), (SM) and (SCO) is strictly more general than assuming (COCO), as we illustrate in Section D.2.
@@
and the $\sqrt{\kappa}$ factor in the denominator is undesirable as it significantly restricts the range of step-size one can use if $\kappa$ is large. Furthermore, in Corollary 3.5 and Theorem 3.6, the factor $q$ becomes $\sqrt{\frac{\ell}{\mu}}=\sqrt{\kappa}$, causing the constant factors in the convergence bounds to potentially become large.
However, there are many cases where $L_{\mathrm{max}}\ll\ell$, showing why it is beneficial to keep the dependency on $L_{\mathrm{max}}$ tight as we do. As an abstract example, when is a generic $\mu$-strongly monotone and $M$-Lipschitz continuous operator, the tight (smallest) cocoercivity parameter one can guarantee on is $\ell=M^{2}/\mu$ [25] (tightness can be shown using, e.g., the scaled relative graph theory in [97], [96, Chapter 13]). On the other hand, we have
@@
which is in contrast with (24) where we used $\ell$ in place of $L_{\mathrm{max}}$ and obtained $\sqrt{\kappa}$ times smaller step-size range. Additionally, note that in this case $q=\frac{L_{\mathrm{max}}}{\sqrt{\ell\mu}}=\frac{L_{\mathrm{max}}}{M}\leq 1$ in Corollary 3.5 and Theorem 3.6, so we can avoid the $\kappa$-dependent factors appearing in the convergence results.
@@
so the Lipschitz constant for $\nabla f_{1}$ with $x^{2}$ fixed (resp. $\nabla f_{2}$ with $x^{1}$ fixed) is $\mu$, i.e., $L_{\mathrm{max}}=\mu$. Therefore, we have $L_{\mathrm{max}}\ll\ell$ in this scenario, as strength of regularization $\mu$ is usually small compared to the smoothness parameter $M$. The same principle applies to the $n$-player analogue of this setup we use in Section 4.2, where each player has the objective function
@@
D.2 Example of non-cocoercive satisfying (CVX), (SM), (QSM) and (SCO)
@@
provided that $N$ is sufficiently large. As a differentiable operator is monotone if and only if $D{\varbbF}+D{}^{\intercal}\succeq 0$ everywhere [96], this shows that is not monotone.

### v2

@@ D.1 Quadratic $n$-player game
D.2 Distributed mobile robot control
We follow the same choice of parameter values $a_{i},b_{i},x^{i}_{\mathrm{anc}},h_{ij}$ within (3) from [KalyvaPsillakis2024_distributed]: $n=5$, $d=1$, $a_{i}=10+i/6$, $b_{i}=i/6$,
@@ D.2 Distributed mobile robot control
$\left(h_{ij}\right)_{\begin{subarray}{c}1\leq i\leq 5\\ 1\leq j\leq 5\end{subarray}}=\begin{bmatrix}0&5&-7&9&-8\\ -5&0&-6&2&-9\\ 7&6&0&7&-4\\ -9&-2&-7&0&-2\\ 8&9&4&2&0\end{bmatrix}$.
We add Gaussian noise with $\sigma^{2}=100$ to the gradients to simulate stochasticity. In this setup, all our theoretical assumptions are satisfied [KalyvaPsillakis2024_distributed].
Appendix E Additional Experiments
E.1 Quadratic $n$-player game with step-size tuning
In this experiment, we simulate the scenario where we do not know the precise theoretical parameters in advance. For each $\tau\in\{1,2,4,5,8,20\}$, we tune $\gamma$ by running PEARL-SGD with each $\gamma\in\{10^{-1},10^{-2},\dots,10^{-6}\}$, and plot the best relative error $\nicefrac{{\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\|^{2}}}{{\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\|^{2}}}$ ($y$-axis) versus the communication round index $p$ ($x$-axis). Figure 5(a) presents results from Deterministic PEARL-SGD, and Figure 5(b) presents results under stochasticity, imposed by mini-batching from the finite sum. The results demonstrate that in practice, we can use $(\tau,\gamma)$ as tunable hyperparameters to achieve the best communication complexity.
(a) Deterministic setup
(b) Stochastic setup
Figure 5: Performance plots for PEARL-SGD on the $n$-player game (2) with different values of $\tau$. For each $\tau$, we use the empirically tuned step-size $\gamma\in\{10^{-1},10^{-2},\dots,10^{-6}\}$ for the best relative error $\frac{\|{\mathbf{x}}_{\tau p}-{\mathbf{x}}_{\star}\|^{2}}{\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\|^{2}}$. Figure 5(a) shows the result from deterministic setup and 5(b) shows the stochastic setup.
E.2 Distributed mobile robot control
Figure 6: Local objective $f_{i}$ in mobile robot control setup.
In Fig.˜6, we plot the local objective values $f_{i}$ in the setup (3) obtained from PEARL-SGD for each robot (player) $i=1,\dots,5$ in the case $\tau=5$. Generally in games, the objectives $f_{i}$ can have both cooperative and competitive components. After the cooperative components in each $f_{i}$ are sufficiently reduced, $f_{i}$’s can oscillate due to the competing interests of players until an equilibrium is found, and then finally stabilize around a certain level.
Appendix F Discussion on Theoretical Assumptions
F.1 Possible simplification of assumptions: Assuming cocoercivity of
@@
In the subsequent paragraphs, we explain in detail why this is the case. However, we emphasize here that if we derived all convergence theory using (COCO) in place of (CVX), (SM) and (SCO) and did not distinguish the role of $L_{i}$’s (the local Lipschitzness parameters from (SM)) from that of $\ell$, then the resulting convergence rates would have become much more pessimistic (worse) in many cases. Therefore, in our work, we choose to use the current set of assumptions. It allows us to more clearly present the tight dependency of convergence rates to $L_{i}$’s. Also note that assuming (CVX), (SM) and (SCO) is strictly more general than assuming (COCO), as we illustrate in Section˜F.2.
@@
and the $\sqrt{\kappa}$ factor in the denominator is undesirable as it significantly restricts the range of step-size one can use if $\kappa$ is large. Furthermore, in Corollary˜3.5 and Theorem˜3.6, the factor $q$ becomes $\sqrt{\frac{\ell}{\mu}}=\sqrt{\kappa}$, causing the constant factors in the convergence bounds to potentially become large.
However, there are many cases where $L_{\mathrm{max}}\ll\ell$, showing why it is beneficial to keep the dependency on $L_{\mathrm{max}}$ tight as we do. As an abstract example, when is a generic $\mu$-strongly monotone and $M$-Lipschitz continuous operator, the tight (smallest) cocoercivity parameter one can guarantee on is $\ell=M^{2}/\mu$ [Facchinei2003FiniteDimensionalVI] (tightness can be shown using, e.g., the scaled relative graph theory in [ryu2022scaled], [ryu2022large, Chapter 13]). On the other hand, we have
@@
which is in contrast with (24) where we used $\ell$ in place of $L_{\mathrm{max}}$ and obtained $\sqrt{\kappa}$ times smaller step-size range. Additionally, note that in this case $q=\frac{L_{\mathrm{max}}}{\sqrt{\ell\mu}}=\frac{L_{\mathrm{max}}}{M}\leq 1$ in Corollary˜3.5 and Theorem˜3.6, so we can avoid the $\kappa$-dependent factors appearing in the convergence results.
@@
so the Lipschitz constant for $\nabla f_{1}$ with $x^{2}$ fixed (resp. $\nabla f_{2}$ with $x^{1}$ fixed) is $\mu$, i.e., $L_{\mathrm{max}}=\mu$. Therefore, we have $L_{\mathrm{max}}\ll\ell$ in this scenario, as strength of regularization $\mu$ is usually small compared to the smoothness parameter $M$. The same principle applies to the $n$-player analogue of this setup we use in Section˜4.1, where each player has the objective function
@@
F.2 Example of non-cocoercive satisfying (CVX), (SM), (QSM) and (SCO)
@@
provided that $N$ is sufficiently large. As a differentiable operator is monotone if and only if $D{\varbbF}+D{}^{\intercal}\succeq 0$ everywhere [ryu2022large], this shows that is not monotone.